A company borrows $75,000 which is to be repaid with equal payments semiannually for 10 years. The interest rate is 10%. Find the semiannual payments.: A company borrows $75,000 which is to be repaid with equal payments semiannually for 10 years. The interest rate is 10%. Find the semiannual payments.
Answer by josmiceli(1851)

Semiannual means twice a years
2*10 = 20payments/ten years
The interest is .1*75000 = 7500
interest/number of payments = interest/payment
7500/20 = 375interest/payment
Let p = payment against principle /payment
(p + 375)*20 = 75000 + 7500
20p + 7500 = 75000 + 7500
20p = 75000
p = 3750
The semiannual payment is
p + 375 = 3750 + 375
p + 375 = 4125
The semiannual payment is $4,125 -answer
check:
4125*20 = 82500
This includes the 10% interest, so subtract it out
82500 - 7500 = 75000
That's the principle